OpenStudy (anonymous):

For the quantum number the =2, how many possible values are there for quantum number m subscript I? A. 2 B. 3 C. 5 D.6

OpenStudy (frostbite):

The magnetic quantum number is ranging from the negative value of the azimuthal quantum number to the positive value.

OpenStudy (anonymous):

So is that mean C?

OpenStudy (anonymous):

C.

OpenStudy (frostbite):

Hmmm nope.

OpenStudy (anonymous):

Are you sure?Cuz i just looked it up and like 3 websites gave me that.

OpenStudy (frostbite):

Lets look at it: \[\Large \sf principal ~ quantum ~ number = n\] \[\Large \sf Azimuthal~quantum~number=0... n-1=l\] \[\Large \sf magnetic~quantum~number=-l... +l=m\] Unless I count totally wrong and need to go to sleep: n=2 highest l = n-1 -> l=2-1 = 1 m = -1 , 0, +1

OpenStudy (anonymous):

So Frostbite, what are you thinking answer wise?

OpenStudy (frostbite):

Should be 3. The values l should be able to take should be -1, 0 and +1 we usually

OpenStudy (frostbite):

Mind I see the web sides @Forever95 ?

OpenStudy (frostbite):

Intresting.

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

Frostbite, you sound like you know what you're doing.. So ill take your word for it.

OpenStudy (anonymous):

its cool.lol

OpenStudy (frostbite):

Did you notice it is for n = 3 on the webside? forever?

OpenStudy (anonymous):

Thank you though @Forever95 :)

OpenStudy (frostbite):

:P

OpenStudy (anonymous):

nope.lol

OpenStudy (anonymous):

blonde moment lol

OpenStudy (anonymous):

@Forever95 I get plenty of those lol

OpenStudy (frostbite):

But I'm glad you think I know what I'm doing since I work with quantum mechanics at the university I study at :P

OpenStudy (anonymous):

yah im just a senior in high school....You would prob get a better chance with his awnser lol

OpenStudy (frostbite):

"Just" is so negative. seniors in high school have lots of knowledge and they still surprise me. Anyway glad we got to a conclusion.

OpenStudy (frostbite):

And welcome to \(\LARGE \sf \bbox[#40B9E9]{\color{white}{Open}}\bbox[#A8CE91]{\color{white}{Study}}\) @winston_salem !

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry had to lol

OpenStudy (anonymous):

@Frostbite I'll let you know if you're correct! :)

OpenStudy (frostbite):

I look forward to hearing I was correct ;)

OpenStudy (frostbite):

But most important: did you understand how I got to the answer?

OpenStudy (anonymous):

3